A buyer can afford no more than $1,200 per month in payments. The most favorable loan available in the market is a 30-year loan at 5%. What is the maximum affordable house with a 20% down payment?

please show using financial calc.
Answer #1

The PV of payments for 30 years
Rate per month = 5%/12
Number of Periods = 20*12 = 360

PV of Payments using financial calculator
N = 360; I/Y = 5%/12; PMT = 1200; CPT PV
PV = 223,537.94

Maximum affordable House = Down payment + PV of Monthly Payments
Maximum affordable House = 20% of Loan amount + 223,537.94

Maximum affordable House = 223,537.94/0.8 = 279,422.43

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well